OpenStudy (megannicole51):

Use integration by parts twice to find the integral of e^(theta)*cos(theta) d(theta)?

OpenStudy (ybarrap):

let u = \(e^\theta\) and \(dv=\cos\theta ~d\theta\) Use this to find \(uv - \int v~du \). You will get $$ e^\theta \sin\theta + \int e^\theta \sin\theta~d\theta $$ Do this again, this time with let u = \(e^\theta\) and \(dv=\sin\theta ~d\theta\). Then you will find that one of the terms will be \(\int e^\theta \cos\theta d\theta\), which is what you started with. Swing this term back over to the other side of the equation, divide by two. That's your answer.

OpenStudy (ybarrap):

Your final answer will be: $$ {e^\theta (\cos\theta+\sin\theta)\over 2} $$

OpenStudy (ybarrap):

Make sense?